OpenStudy (anonymous):

q + 12 – 2(q – 22) > 0? Help please. I don't understand..

OpenStudy (anonymous):

q + 12 - 2 (q-22) > 0 You need the simplify the LHS, make it: q + 12 - (2q-44) = q + 12 - 2q + 44 = -q + 56, which you want as positive. So, we have: \[ -q + 56 > 0 \rightarrow \boxed{q<56}. \]Hope that helped! Let us know if you still have a question. We'd be very glad to help. :D